# Meta-analysis of median survival times with inverse-variance weighting

### v1

@@ Abstract
We consider the problem of meta-analyzing outcome measures based on median survival times, such as the difference of median survival times between groups. Primary studies with time-to-event outcomes often report estimates of median survival times and corresponding confidence intervals based on the Kaplan-Meier estimator. However, applying conventional inverse-variance weighted methods to meta-analyze outcome measures based on median survival is often challenging because within-study standard error estimates are typically not available in this setting. In this article, we consider an inverse-variance weighted approach to meta-analyze median survival times that estimates the within-study standard errors from the reported confidence intervals. We conduct a series of simulation studies evaluating the performance of this approach at the study level (i.e., for estimating the standard error of median survival) and the meta-analytic level (i.e., for estimating the pooled median, difference of medians, and ratio of medians). We find that this approach often performs comparable to a benchmark approach that uses the true within-study standard errors for meta-analyzing median-based outcome measures. We then illustrate an application of this approach in a meta-analysis evaluating survival benefits of being assigned to experimental arms versus comparator arms in randomized trials for non-small cell lung cancer therapies.
Keywords— meta-analysis, median, survival outcomes, time-to-event outcomes, aggregate data; inverse-variance weighting
@@ 1 Introduction
Researchers and policy-makers are often tasked with synthesizing survival (time-to-event) data across studies in meta-analyses. For example, decisions about drug reimbursement or clinical practice guidelines may be based on such analyses. When meta-analyses of survival data are performed, the hazard ratio is the typical outcome measure used [1]. Standard inverse-variance weighted approaches are typically applied to meta-analyze hazard ratios, where study-specific estimates of hazard ratios and their standard errors are often reported or can be estimated from various sets of summary statistics extracted from the primary studies [2].
@@
However, the hazard ratio may not be the most appropriate outcome measure in some time-to-event studies and meta-analyses of them. Many authors argue that the interpretation of hazard ratios is not straightforward, especially when hazard ratios vary over time (i.e., the assumption of proportional hazards is violated) [3, 4, 5, 6]. Furthermore, in causal inference settings, (counterfactual) hazard ratios do not have a clear causal interpretation [7, 8]. In such cases, conclusions from analyses based on hazard ratios can be misleading. Substantive interest may instead lie in quantifying how typical survival times differ between groups, such as the difference of median survival times between groups.
@@
We consider the problem of meta-analyzing outcome measures based on median survival times. Primary studies with time-to-event outcomes often report estimates of median survival times and corresponding confidence intervals due to their interpretability. For example, in a meta-analysis performed by Iskander et al. [9], 94% (120 out of 128) of primary studies reported estimates of median progression-free survival times and confidence intervals separately for cancer patients randomized to experimental arms and comparator arms. In many cases, the full survival curves are not available in the primary studies. For example, this may happen when the outcome of interest in the meta-analysis is a secondary outcome in the primary study. It may also occur when data are extracted from sources such as ClinicalTrials.gov, where survival curves are not reported and full publications may not be available (e.g., [9]). Although outcome measures based on median survival may be of interest and data on median survival are often available, meta-analyzing median-based outcome measures can be challenging.
@@
A key challenge for meta-analyzing an outcome measure based on median survival times is that standard inverse-variance weighted meta-analytic methods cannot necessarily be directly applied without bias. Standard meta-analytic methods require an estimate of the standard error of the outcome measure estimate in each primary study. Primary studies rarely report estimates of the standard error of the median survival time estimate. Rather, studies typically report nonparametric confidence intervals around median survival time estimates based on inverting Kaplan-Meier estimates of the survival function [10]. Unlike confidence intervals for hazard ratios [1], confidence intervals for median survival are usually not constructed based on the Wald method and consequently standard error estimates cannot necessarily be back-computed without bias. Poor estimation of within-study standard errors can potentially have several negative downstream consequences in meta-analysis, including (i) reducing the efficiency of the pooled outcome estimator, (ii) poor estimation of the between-study heterogeneity in random effects meta-analyses, and (iii) poor confidence interval coverage for the pooled outcome in common effect meta-analyses (e.g., [11]).
@@
In this work, we evaluate the performance of an inverse-variance weighted approach to meta-analyzing median survival times that is based on treating confidence intervals around median survival time estimates as Wald-type intervals. We refer to this approach as the Wald approximation-based approach. We consider the setting where the meta-analysis consists of one-group studies (i.e., where the target of inference is the median survival time) as well as two-group studies (i.e., where the target of inference is the difference or ratio of median survival times between groups). In simulation studies, we systematically evaluate the performance of the Wald approximation-based approach and compare it to a benchmark approach that uses the true (unknown) standard error of the medians. Finally, we illustrate an application of the Wald approximation-based approach in a meta-analysis of median overall survival times of clinical trial participants with non-small cell lung cancer.
@@ 2.1 Standard meta-analytic methods
where the are considered to be known quantities. In the setting of a common effect meta-analysis, it is assumed that the true study-specific outcome measures, , are identical, which we denote by . In the setting of a random effect meta-analysis, the true study-specific outcome measures are assumed to be distributed as
@@ 2.2.1 Study level
where denotes the quantile of the standard normal distribution and denotes an estimate of the standard error of . For example, may be the conventional estimator given by
where is based on the Greenwood formula and is an estimator of the probability density function for the event times [16] (see Appendix A for more details). Following (1) and (2), we can back-compute by
@@
As discussed in Section 1, in practice the confidence interval may not be a Wald interval and the particular method used to construct the interval is often unknown to the meta-analyst. In Remark 1, we provide some intuition on why the Wald approximation in (3) may perform well for commonly used (non-Wald) confidence intervals in this context. Note that we systematically evaluate the performance of the Wald approximation for commonly used confidence intervals in Section 3.
@@
There is an alternative interpretation of the standard error estimator in (3) that is based on other assumptions on how the confidence interval was constructed. First, consider the setting where there is no censoring. Suppose that and are central order statistics such that and . For example, the confidence interval based on inverting a two-sided sign test for the median is one such interval. Under mild regularity conditions, Ozturk and Balakrishnan [17] showed that (3) is a consistent estimator of based on a Bahadur-type representation of central order statistics. In fact, Ozturk and Balakrishnan [17] and Lang et al. [18] used this standard error estimator to meta-analyze medians and differences of medians with inverse-variance weighting in non-survival settings. In survival settings with right-censoring, a similar justification for the standard error estimator of (3) can be made based on a Bahadur-type representation of the Kaplan-Meier estimator of quantiles of the event time distribution [19]. In fact, the commonly used Brookmeyer-Crowley method [10] for constructing confidence intervals around median survival can be seen as inverting a generalization of the sign test for the median in settings with right-censoring, which suggests the estimator in (3). Despite this alternative interpretation, we will refer to the estimator in (3) as the Wald approximation-based approach for simplicity.
@@
For simplicity, we motivated the Wald approximation based on the setting of using the Wald method without a transformation and with equal tail probabilities (i.e., ). It is straightforward to see via Taylor expansions that (3) can also be motivated in scenarios where the confidence interval was constructed on a different scale (e.g., log scale) and was back-transformed to the original scale. Moreover, when the confidence interval has unequal tail probabilities, we illustrate in Appendix B that the Wald approximation will overestimate the standard error although the degree of bias is small unless one of the tail probabilities is nearly 0.
@@ 2.2.2 Meta-analytic level
To meta-analyze any of the outcome measures, we first estimate the standard error of the median estimator (in each group) as described in the previous subsection. One can immediately proceed to meta-analyze the median in a single group. To meta-analyze the difference of medians, we need to estimate the standard error of the difference of medians estimator. Due to the independence of the two groups, we can estimate we estimate the standard error by
@@
After obtaining a pooled estimate of the log ratio of medians, the inverse transformation is applied. That is, the pooled estimate of the ratio of medians is .
@@ 3 Simulations
In this section, we present simulation studies evaluating the performance of the Wald approximation-based approach at both the study level and meta-analytic level. Specifically, we evaluate the performance of the Wald approximation-based approach for estimating the standard error of median survival estimates in Section 3.1. We evaluate the Wald approximation-based approach for meta-analyzing the median, difference of medians, and ratio of median survival times in Section 3.2.
@@ 3.1.1 Data generation
To generate each longitudinal data set, we simulated underlying event times () from either (i) an exponential distribution with a mean of 40, (ii) a Weibull distribution with a shape parameter of and a scale parameter of , or (iii) a two-component mixture of Weibull distributions. The first component had a probability of with parameters and the second component had a probability of with parameters . Plots of the density functions of these distributions are given in Appendix C.1. Individuals were considered administratively lost to follow-up if their underlying event times exceeded 100. The censoring times were independently simulated from either a distribution or an exponential distribution with a mean of 60. We observe the minimum of the underlying event time and the censoring time. The percentage of individuals censored for each combination of event time distribution and censoring time distribution is listed in Appendix C.1.
@@
We applied the Wald approximation to estimate the standard error of the median survival estimate for each type of confidence interval in each longitudinal data set. We then compared the estimated standard errors to the true standard errors of the median estimator. Specifically, we computed the relative bias of the standard error estimator. Letting denote the estimated standard error of the median estimator () in the th iteration and denote the true value of the standard error, the relative bias is defined as the average of across the iterations. The true standard errors were obtained by performing Monte Carlo integration with samples.
@@ 3.1.2 Results
Figure 1 illustrates the distributions of the estimated standard errors in the scenarios with the exponential and Weibull distributions. The results for the mixture distribution are placed in Appendix C.1, as similar trends held. Appendix C.1 also lists the relative bias values in all scenarios.
@@ Results
Similar to the simulations for meta-analyzing median survival, we found that the Wald approximation-based approach performed very similarly to the benchmark approach. Both methods had nearly identical values for the bias, standard error, and coverage of their 95% confidence intervals for the pooled difference of medians and the between-study variance in each scenario. Their confidence intervals for the pooled difference of medians and between-study variance had nominal or near-nominal coverage in all scenarios.

### v2

@@ Abstract
We consider the problem of meta-analyzing outcome measures based on median survival times. Primary studies with time-to-event outcomes often report estimates of median survival times and confidence intervals based on the Kaplan-Meier estimator. However, outcome measures based on median survival are rarely meta-analyzed, as standard inverse-variance weighted methods require within-study standard errors that are typically not reported. In this article, we consider an inverse-variance weighted approach to meta-analyze median survival times that estimates the within-study standard errors from the reported confidence intervals. We show that this method consistently estimates the standard error of median survival when applied to confidence intervals constructed by the Brookmeyer-Crowley method. We conduct a series of simulation studies evaluating the performance of this approach at the study level (i.e., for estimating the standard error of median survival) and the meta-analytic level (i.e., for estimating the pooled median, difference of medians, and ratio of medians) for commonly used confidence intervals for median survival, including the Brookmeyer-Crowley method and nonparametric bootstrap. We find that this approach often performs comparably to a benchmark approach that uses the true within-study standard errors for meta-analyzing median-based outcome measures when within-study sample sizes are moderately large (e.g., above 50). However, when the effective sample sizes are small, the method can yield biased estimates of within-study standard errors. We illustrate an application of this approach in a meta-analysis evaluating survival benefits of being assigned to experimental arms versus comparator arms in randomized trials for non-small cell lung cancer therapies.
Keywords— meta-analysis, median survival, time-to-event outcomes, aggregate data, inverse-variance weighting
@@ 1 Introduction
Researchers and policy-makers are often tasked with synthesizing time-to-event (survival) data across studies in meta-analyses. For example, decisions about drug reimbursement or clinical practice guidelines may be based on such analyses. When meta-analyses of time-to-event data are performed, the hazard ratio is the typical outcome measure used [1]. Standard inverse-variance weighted approaches are typically applied to meta-analyze hazard ratios, where study-specific estimates of hazard ratios and their standard errors are often reported or can be estimated from various sets of summary statistics extracted from the primary studies [2].
@@
However, the hazard ratio is not the most appropriate outcome measure in some time-to-event studies and meta-analyses of them. Many authors argue that the interpretation of hazard ratios is not straightforward, especially when hazard ratios vary over time (i.e., the assumption of proportional hazards is violated) [3, 4, 5, 6]. Furthermore, in causal inference settings, (counterfactual) hazard ratios do not have a clear causal interpretation [7, 8]. In such cases, conclusions from analyses based on hazard ratios can be misleading. Substantive interest may instead lie in quantifying how typical survival times differ between groups, such as the difference of median survival times between groups.
@@
We consider the problem of meta-analyzing outcome measures based on median survival times. Primary studies with time-to-event outcomes often report estimates of median survival times and corresponding confidence intervals due to their interpretability. For example, in a meta-analysis performed by Iskander et al. [9], 94% (120 out of 128) of primary studies reported estimates of median progression-free survival times and confidence intervals separately for cancer patients randomized to experimental arms and comparator arms. In many cases, the full survival curves are not available in the primary studies. For example, this may happen when the outcome of interest in the meta-analysis is a secondary outcome in the primary study. It may also occur when data are extracted from sources such as ClinicalTrials.gov, where survival curves are not reported and full publications may not be available (e.g., [9]).
@@
Although outcome measures based on median survival are often of interest and median survival times are routinely reported in time-to-event studies, meta-analyses of median-based outcome measures are relatively uncommon, likely due to methodological challenges that are understudied (see also Section 5.1). For example, the Cochrane Handbook for Systematic Reviews of Interventions [1] does not discuss methods to meta-analyze outcome measures based on median survival. A key challenge for meta-analyzing an outcome measure based on median survival times is that primary studies rarely report estimates of the standard error of the median survival time estimate, which is needed for applying standard inverse-variance weighted meta-analytic methods. Rather, studies typically report nonparametric confidence intervals around median survival time estimates based on inverting Kaplan-Meier estimates of the survival function [10]. Unlike confidence intervals for hazard ratios [1], confidence intervals for median survival are usually not constructed based on the Wald method. Consequently, the validity of back-computing standard error estimates using the standard Wald formula is not guaranteed. Poor estimation of within-study standard errors can potentially have several negative downstream consequences in meta-analysis, including (i) reducing the efficiency of the pooled outcome estimator, (ii) poor estimation of the between-study heterogeneity in random effects meta-analyses, and (iii) poor confidence interval coverage for the pooled outcome in common effect meta-analyses (e.g., [11]).
@@
In this work, we investigate the validity of an inverse-variance weighted approach to meta-analyzing median survival times that is based on treating confidence intervals around median survival time estimates as Wald-type intervals. We refer to this approach as the Wald approximation-based approach. We show that the Wald approximation yields a consistent estimator of the standard error of the median survival estimate for the commonly used Brookmeyer-Crowley confidence interval method [10]. In simulation studies, we evaluate the finite sample performance of the Wald approximation for estimating standard errors based on different types of confidence intervals. We then investigate the application of the Wald approximation in meta-analysis, where we consider the setting where the meta-analysis consists of one-group studies (i.e., where the target of inference is the median survival time) as well as two-group studies (i.e., where the target of inference is the difference or ratio of median survival times between groups). In additional simulation studies, we evaluate the performance of the Wald approximation-based approach in meta-analysis and compare it to a benchmark approach that uses the true (unknown) standard errors of the medians. Finally, we illustrate an application of the Wald approximation-based approach in a meta-analysis of median overall survival times of clinical trial participants with non-small cell lung cancer.
@@ 2.1 Standard meta-analytic methods
where the are considered to be known quantities. In the setting of a common effect meta-analysis, it is assumed that the true study-specific outcome measures, , are identical, which we denote by . In the setting of a random effects meta-analysis, the true study-specific outcome measures are assumed to be distributed as
@@ 2.2.1 Study level
where denotes the quantile of the standard normal distribution and denotes an estimate of the standard error of . For example, may be the conventional estimator given by where is based on the Greenwood formula, is the sample size, and is an estimator of the probability density function for the event times [16] (see Section 1 of the Supplementary Material for more details). Following (1) and (2), we can back-compute by
@@
For simplicity, we motivated the Wald approximation based on the setting of using the Wald method without a transformation and with equal tail probabilities (i.e., ). It is straightforward to see via Taylor expansions that (3) can also be motivated in scenarios where the confidence interval was constructed on a different scale (e.g., log scale) and was back-transformed to the original scale. Moreover, when the confidence interval has unequal tail probabilities, we illustrate in Section 2 of the Supplementary Material that the Wald approximation will overestimate the standard error although the degree of bias is small unless one of the tail probabilities is nearly 0.
@@
An upper limit for the confidence interval for may not be reported. In this case, if the confidence interval was constructed using the Wald method, one can estimate the standard error by .
Brookmeyer-Crowley intervals
As discussed in Section 1, confidence intervals for median survival reported in practice are typically not Wald intervals. Rather, such confidence intervals are commonly constructed using the test-inversion method proposed by Brookmeyer and Crowley [10]. Unlike Wald intervals, Brookmeyer-Crowley intervals are not symmetric around the median estimate, which raises questions about whether the Wald approximation in (3) remains valid. Here, we show that the Wald approximation yields a consistent estimator of the standard error of the Kaplan-Meier estimator of median survival under this confidence interval method.
We first introduce the following notation and assumptions. For simplicity, we drop the study () and group () subscripts in this subsection. Let denote i.i.d. event times with distribution function , density , and survival function . Let denote i.i.d. censoring times that are independent of the event times. The observed data consist of and for . The population median survival time is given by . We assume that is continuous and positive at (so ), and that for some to ensure identifiability and validity of the asymptotic theory.
Let denote the Kaplan-Meier estimator of the survival function, and let denote the Kaplan-Meier estimator of the median. Let denote the asymptotic variance of , and let denote its estimator based on the Greenwood formula (see Section 1 of the Supplementary Material for details). The Brookmeyer-Crowley method constructs a confidence interval for the median by inverting a hypothesis test. Specifically, the confidence interval is defined as
with and denoting its infimum and supremum, respectively. We use uppercase to denote the confidence interval limits as random variables (emphasizing their dependence on sample size ), in contrast to the lowercase used elsewhere to denote observed confidence limits extracted from studies. This interval may be constructed on the original scale (as shown above) or on a transformed scale (e.g., using log or log-minus-log transformations of the survival function) and then back-transformed.
Under these conditions, the Kaplan-Meier estimator of the median is asymptotically normally distributed:
where denotes convergence in distribution as and . The following result establishes that the Wald approximation (scaled by ) provides a consistent estimator of the asymptotic standard error, .
Proposition 1.
Under the given assumptions, the Wald approximation for the standard error satisfies:
where denotes convergence in probability as .
The proof is provided in Section 3 of the Supplementary Material. The proof uses the stochastic equicontinuity of the Kaplan-Meier process and Taylor expansions of the survival function around . The key insight is that when expressing the interval width , the empirical process terms appearing in and cancel out, leaving only a deterministic term involving and asymptotically negligible remainders. A similar result was shown by Ozturk and Balakrishnan [17] in non-survival settings (i.e., without censoring) based on Bahadur-type representations of sample quantiles.
Despite this alternative interpretation, we will refer to the estimator in (3) as the Wald approximation-based approach for simplicity.
@@ 2.2.2 Meta-analytic level
To meta-analyze any of the outcome measures, we first estimate the standard error of the median estimator (in each group) as described in the previous subsection. One can immediately proceed to meta-analyze the median in a single group. To meta-analyze the difference of medians, we need to estimate the standard error of the difference of medians estimator. Due to the independence of the two groups, we can estimate the standard error by
@@
After obtaining an estimate of the pooled log ratio of medians, the inverse transformation is applied. That is, the estimate of the pooled ratio of medians is . We elaborate on the interpretation of the pooled ratio of medians in Remark 3.
@@ 3 Simulations
In this section, we present simulation studies evaluating the performance of the Wald approximation-based approach at both the study level and meta-analytic level. Specifically, we evaluate the performance of the Wald approximation-based approach for estimating the standard error of median survival estimates in Section 3.1. We evaluate the Wald approximation-based approach for meta-analyzing the median, difference of medians, and ratio of median survival times in Section 3.2. The code for reproducing these simulation studies is available at https://github.com/stmcg/ma-median-survival.
@@ 3.1.1 Data generation
To generate each longitudinal data set, we simulated underlying event times () from either (i) an exponential distribution with a mean of 40, (ii) a Weibull distribution with a shape parameter of and a scale parameter of , or (iii) a two-component mixture of Weibull distributions. The first component had a probability of with parameters and the second component had a probability of with parameters . Plots of the density functions of these distributions are given in Section 4.1 of the Supplementary Materials. Individuals were considered administratively lost to follow-up if their underlying event times exceeded 100. The censoring times were independently simulated from either a distribution or an exponential distribution with a mean of 60. We observe the minimum of the underlying event time and the censoring time. The percentage of individuals censored for each combination of event time distribution and censoring time distribution is listed in Section 4.1 of the Supplementary Materials. The censoring rates ranged between 26% and 41%.
@@
We applied the Wald approximation to estimate the standard error of the median survival estimate for each type of confidence interval in each longitudinal data set. We then compared the estimated standard errors to the true standard errors of the median estimator. Specifically, we computed the relative bias of the standard error estimator. Letting denote the estimated standard error of the median estimator () in the th iteration and denote the true value of the standard error, the relative bias is defined as the average of across the iterations. The true standard errors were obtained by performing Monte Carlo integration with samples (see Section 4.6 of the Supplementary Material for more details).
@@ 3.1.2 Results
Figure 1 illustrates the distributions of the estimated standard errors in the scenarios with the exponential and Weibull distributions. The corresponding figure for the mixture distribution is placed in Section 4.1 of the Supplementary Material, as similar trends held. Section 4.1 of the Supplementary Material also lists the relative bias values in all scenarios.
@@
3.1.3 Additional study-level simulations
We performed a series of additional study-level simulation studies to explore the sensitivity of the Wald approximation to (i) small sample sizes, (ii) a higher degree of censoring, (iii) extremely skewed event time distributions, and (iv) the number of bootstrap replicates when bootstrap confidence intervals are reported.
Impact of smaller sample sizes
To evaluate the performance of the Wald approximation in settings where the asymptotic normality of the Kaplan-Meier estimator may not serve as a good approximation, we performed additional study-level simulations with smaller sample sizes. Specifically, we considered sample sizes of while keeping the same event time and censoring distributions as described in Section 3.1.1.
The results for the scenarios with the exponential event time distribution are given in Figure 2, and the corresponding figure for the Weibull distribution and mixture distributions is placed in Section 4.2 of the Supplementary Material. The Wald approximation often exhibited considerable bias when . For example, the relative bias was up to -39% in one scenario (exponential event time distribution, exponential censoring distribution). In the scenarios with , the bias was significantly reduced, yielding results comparable to the scenarios where .
Impact of a higher degree of censoring
We also investigated the impact of having a higher degree of censoring on the performance of the Wald approximation. We modified the data generation mechanism described in Section 3.1.1 to include higher degrees of censoring by generating censoring times from exponential distributions with means of 40 and 25. We considered the same event time distributions and sample sizes as in Section 3.1.1. These scenarios had censoring rates ranging between 43% and 65%.
The results for these scenarios are placed in Section 4.3 of the Supplementary Materials for parsimony. The Wald approximation performed poorly in scenarios with the highest degree of censoring and smallest sample sizes. For example, the relative bias was -39% in the scenario with high censoring and . However, even in the settings with the highest censoring, the Wald approximation performed well when the sample size was large. For example, the relative bias was below 2% in all scenarios with .
Impact of highly skewed event time distributions
We additionally performed simulations with very highly skewed event time distributions in Section 4.4 of the Supplementary Material. When the sample size was small, the Wald approximation clearly showed bias, especially when using the Brookmeyer-Crowley method based on the log transformation. In the most extreme case, a relative bias of 19% was observed. However, the bias generally decreased as the sample size increased. When the sample size was 1000, the relative bias was below 2% in all these scenarios as well.
Impact of the number of bootstrap replicates
We investigated sensitivity of the Wald approximation to the number of bootstrap replicates when nonparametric bootstrap confidence intervals are reported. The results are presented in Section 4.5 of the Supplementary Material. Increasing the number of bootstrap replicates did not meaningfully affect the bias or precision of the Wald approximation.
@@ Results
Similar to the simulations for meta-analyzing median survival, we found that the Wald approximation-based approach performed very similarly to the benchmark approach. Both methods had nearly identical values for the bias, standard error, and coverage of their 95% confidence intervals for the pooled difference of medians and the between-study variance in each scenario. Their confidence intervals for the pooled difference of medians and between-study variance had nominal coverage or slight overcoverage in all scenarios. The slight overcoverage of the confidence interval for the pooled difference of medians may be due to using Hartung-Knapp adjustment in the random effects scenarios and t-interval in the common effect scenario.
@@
3.2.4 Impact of misspecifying the form of between-study heterogeneity
The meta-analysis of the difference of medians and the meta-analysis of the ratio of medians methods involve different assumptions on the form of the between-study heterogeneity. The difference of medians meta-analysis assumes heterogeneity on the additive scale (i.e., the true study-specific differences of medians are normally distributed) and the ratio of medians meta-analysis assumes heterogeneity on the multiplicative scale (i.e., the true study-specific log ratios of medians are normally distributed). Consequently, when applying both of these methods to the same data (e.g., as in Section 4), one may be concerned about misspecifying the form of between-study heterogeneity in at least one of these methods. In Section 5.1 of the Supplementary Material, we evaluate the performance of these two methods in our simulation study when misspecifying the form of the between-study heterogeneity in this manner. We found that such misspecification had little impact on the performance of these methods. Related, see [26] for a discussion of implications of the distributional assumptions on between-study heterogeneity.
3.2.5 Impact of small within-study sample sizes
We additionally performed analogous simulations where the within-study sample sizes were considerably smaller. As the study-level simulations found that the Wald approximation can perform poorly when sample sizes are small, these simulations illustrate the downstream consequences of applying the Wald approximation in such meta-analyses. Specifically, we re-performed the simulation studies of the meta-analysis of median survival (Section 3.2.1), difference of median survival (Section 3.2.2), and ratio of median survival (Section 3.2.3), where the sample sizes were drawn from a discrete distribution rather than a discrete distribution. Detailed results are given in Section 5.2 of the Supplementary Material.
For all three simulation studies, the Wald approximation-based approach often remained unbiased for estimating the pooled outcome measure and generally had similar efficiency as the benchmark approach. However, the Wald approximation-based approach had below nominal coverage of its 95% confidence intervals for the pooled outcome measure in the common effect scenarios (e.g., dropping to 0.65 for the pooled median). The Wald approximation-based approach also generally overestimated the between-study variance and had below nominal coverage for its 95% confidence intervals for this target parameter. Both of these issues can be seen as consequences of the Wald approximation systematically underestimating the within-study variances, as found in the study-level simulations with small sample sizes.